XYY manufactures fibre optic cables. This is an increasingly competitive market and the Chief Financial Officer (CFO) of Venus has suggested that the company’s pricing policy should change. Specifically, he has suggested that Venus’ quotation for a new contract should be based on target costing. The costs relevant to this are listed below and the prices in this market are all quoted on a per metre basis:

Target price   $196 per metre of cable.

The contract is estimated at 25,000 metres of cable.

Venus profit margin     18%.

Material cost $50 per metre.

Production of each metre takes 0.5 hours of labour at a cost of $30 per hour.

Each metre of cable will use 1.2 units of machine time at a cost of $43 per hour.

Delivery and installation will cost $12 per metre of cable.

Total design costs for all 25,000 metres of cable are expected to be $900,000.

The rate of failed metres of cable is expected to amount to 3%. These metres will be reworked at an average cost of $35 per metre.

Required:

  1. Calculate the target cost for the new contract and the cost gap.

b.Comment on actions that Venus might take to reduce the cost gap.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
XYY
Answer a Amount $ Amount $ Note
Target price                     196.00 A
Profit margin 18% B
Target profit                       35.28 C=A*B
Target cost per meter                     160.72 D=A-C
Contract (meters)               25,000.00 E
Target cost for 25,000 meters         4,018,000.00 F=D*E
Calculation of actual cost of 25,000 meters
Material cost per meter                 50.00 G
Contract (meters)         25,000.00 See E
Material cost for 25,000 meters         1,250,000.00 H=G*E
Labor hours for each meter                   0.50 I
Contract (meters)         25,000.00 See E
Labor hours for 25,000 meters         12,500.00 J=I*E
Labor rate per hour                 30.00 K
Labor cost for 25,000 meters             375,000.00 L=J*K
Machine hours for each meter                   1.20 M
Contract (meters)         25,000.00 See E
Machine hours for 25,000 meters         30,000.00 N=M*E
Machine rate per hour                 43.00 O
Machine cost for 25,000 meters         1,290,000.00 P=N*O
Delivery and installation cost per meter                 12.00 Q
Contract (meters)         25,000.00 See E
Delivery and installation cost for 25,000 meters             300,000.00 R=Q*E
Total design costs for 25,000 meters             900,000.00 S
Failure rate 3% T
Contract (meters)         25,000.00 See E
Failure meters               750.00 U=T*E
Rework cost per meter                 35.00 V
Rework cost for failed cable meters               26,250.00 W=U*V
Actual cost for 25,000 cable meters         4,141,250.00 X= H+L+P+R+S+W
COST GAP
Actual cost for 25,000 cable meters          4,141,250.00 See X
Less: Target cost for 25,000 meters          4,018,000.00 See F
COST GAP             123,250.00 Y=X-F
Answer b
There is a cost gap of $ 123,250 for 25,000 meters of cable.
It can be reduced if Venus could take following steps:
1. Try to keep failure at the lowest. Failed items are nothing but quality control issues. It contributed $ 26,250 of the total cost gap. It should increase supervision to keep the failure rates to the minimum.
2. Review the supply chain management. It should look for those suppliers who can sell the raw materials at a convenient price without affecting the quality. It should not rely on 1 supplier rather approach five suppliers and review their quotes to get the lowest priced quality materials.
3. Delivery and installation are part of the sell process but it does not add value to the production. It should review this cost model and to arrive at more cost efficient options.
4. Total design costs are allocated to only 25,000 meters. If the company has option to sell more than 25,000 meters then it will automatically reduce the design costs per meter.